Unless stated otherwise, interest is compounded annually, and payments occur at the end of the period. Face value for bonds is $1000.

  1. Tower preferred stock has a $50 face value and an annual preferred dividend rate of 6.2%. If the price is $47, find the required return.
Answer #1

Annual dividend=50*6.2%=3.1

Required rate of return=Annual dividend/Current value

=3.1/47

=6.60%(Approx).